OpenStudy (anonymous):

A scientist has discovered an organism that replicates at a rate of five per hour and lives for one week. Each replicated organism also replicates at the same rate. At hour one, there is one organism. At hour two, there are five more organisms. How many total organisms are there at hour seven? 2,801 19,531 19,607 97,655

OpenStudy (h2so4):

Is it 97,655?

OpenStudy (cuanchi):

how did you got that?

OpenStudy (aaronq):

I'm probably not seeing something because I got none of those answers.. We know that after each hour each old organism gives rise to 5 new organisms. Hr 1, 1 organism Hr 2, 5 new organisms + 1 old organism = 6 ^ (1*5) Hr 3, 30 new + 6 old =36 ^ (6*5) Hr 4, 180 new + 36 old= 216 ^ (36*5) If we continue this... Hr 1 2 3 4 5 6 7 # of Organisms 1 6 36 216 1296 7776 46656

OpenStudy (cuanchi):

Yes I got the same that @aaronq , at least we are missing something here